Solve by Factoring
12x(Squared) +19x-18= 0

Found 2 solutions by scott8148, nerdybill:
Answer by scott8148(6628)

factors of 12 are 3 and 4
factors of -18 are -2 and 9

(3x-2)(4x+9)=0

3x-2=0 ___ 3x=2 ___ x=2/3

4x+9=0 ___ 4x=-9 ___ x=-9/4

Answer by nerdybill(7384)

12x(Squared) +19x-18= 0
(3x-2)(4x+9) = 0
.
Setting each to zero:
(3x-2)= 0
3x = 2
x = 2/3
and
(4x+9) = 0
4x = -9
x = -9/4
.
x = {-9/4, 2/3}